solve for x
1/2x-2=1/3(x-2)

Answer by checkley71(8403)

1/2x-2=1/3(x-2)
1/2x-2=x/3-2/3
x/2-x/3=2-2/3
(3x-2x)/6=1 1/3
x/6=4/3 cross multiply
3x=24
x=24/3
x=8 answer.
proof
1/2*8-2=1/3(8-2)
4-2=1/3*6
2=2